OpenStudy (anonymous):

A ship travels 50 km on a bearing of 27 degrees, then travels on a bearing of 117 degrees for 140 km. Find the distance traveled from the starting point ot the ending point.

hero (hero):



hero (hero):

That's a rough sketch. For more accuracy, use graph paper and protractor.

OpenStudy (wolf1728):

Hero - I'm working on this too and instead of graphing a solution I'm using trig

hero (hero):

You should always sketch when there's an opportunity to. It's one of the first steps in math problem solving. It provides a solid basis of approach when solving a problem. The point of solving math is to take something that is by its own nature abstract and intangible and covert to something that is tangible, concrete, and visual for demonstration purposes.

hero (hero):

Are you wolfgirl?

OpenStudy (wolf1728):

No I am not wolfgirl

OpenStudy (wolf1728):

The figures I am going to post are from the attached graphic. The northern position = opposite + opp2 = 22.6995 + 44.5505 which equals 67.25 km North of the start position For the western position, we have to subtract the adj position from the adj2 position. So western position is adj2 - adj = 63.5577 - 44.5505 = 19.0072 Western position is 19.0072 km west of the starting position.